i REALLY dont get this whole concept, but here's the question!
You randomly choose a number from 0 to 7. What are the odds that the number is four or more?
Thanks!

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

You randomly choose a number from 0 to 7. What are the odds that the number is four or more?
---------------
4 or more is 4, 5, 6 & 7, 4 possibilities
The total # of possibilities is 8.
4 of 8 = 1/2
odds are 1 out of 2, or 1 to 1
--------
"odds" is not a clearly defined term.

Answer by stanbon(75887)

You randomly choose a number from 0 to 7. What are the odds that the number is four or more?
-----
P(4 or more) = 4/8 = 1/2
----
P(NOT 4 or more) = 4/8 = 1/2
-----
odds FOR 4 or more = [P(4 or more)]/[P(NOT 4 or more)] = [1/2] / [1/2)] = 1:1
================================================================================
Cheers,
Stan H.